# Traversable wormholes inside anisotropic magnetized neutron stars: physical properties and potential observational imprints

### v1

@@ Abstract
In this paper, we formulate wormhole-plus-neutron-star (WH+NS) systems supported by two scalar fields, allowing for both pressure anisotropy and magnetic fields. In general, such WH+NS systems contain ghosts; however, these ghosts can be eliminated. We find that neither anisotropy nor magnetic fields affect the traversability of the wormhole. In particular, the null energy condition (NEC) remains violated in the vicinity of the wormhole throat, ensuring the traversable nature of the geometry. For magnetized configurations, the resulting WH+NS systems can become extremely massive, with ADM masses exceeding $8\,M_{\odot}$, and can exhibit large surface redshifts exceeding $z\simeq 1.5$. Furthermore, we analyze the gravitational-wave echo time of the systems, which serves as a potential observational imprint. Our results indicate that the echo time can vary depending on the magnetic field configuration, suggesting that WH+NS systems may provide distinctive signals of gravitational echo.
preprint: APS/123-QED
I Introduction
@@
Wormholes are intriguing constructs that serve as tunnels connecting distant spacetime points [1, 2, 3, 4, 5, 6, 7, 8]. Both distinct spacetime domains are linked through a throat-like geometry. Originally, wormholes emerged as exact solutions of the Einstein field equation [9]. Due to their geometric nature, these exotic objects may potentially exist in curved spacetimes within a broad class of gravity theories [10]. Interestingly, wormholes have been associated with rapid interstellar travel, warp drives [11], and candidates for active galactic nuclei [12].
@@
Historically, the first insight into what would later be interpreted as a wormhole can be traced back to Flamm’s 1916 work [13] (see Ref. [14] for the English translation), where he analyzed the Schwarzschild spacetime. In 1935, Einstein & Rosen [15] found a solution bearing a resemblance to a wormhole, which was identified as a “bridge” that connects two identical physical spaces. This work gave rise to the term “Einstein–Rosen bridge”. The term “wormhole” was later introduced in 1957 by Misner & Wheeler [16]. It is worth noting that the putative wormhole in the Einstein–Rosen bridge is not traversable, since it is colocated with the blackhole’s singularity. The surface of the blackhole is the event horizon, which cannot be a wormhole [17].
@@
The formalism of traversable wormholes was introduced by Morris & Thorne [18] in their seminal 1988 paper. In constructing traversable wormholes, the null energy condition (NEC) must be violated by the energy-momentum tensor (EMT) [19]. It should be noted that there are three other energy conditions: the weak energy condition (WEC), the strong energy condition (SEC), and the dominant energy condition (DEC) [17]. The violation of NEC implies that all the energy conditions are not validated [20]. In general relativity (GR), the violation of NEC typically requires the presence of exotic matter [21]. Many authors [22, 23, 24, 25, 26, 27] have included the exotic matter in their studies on wormholes. Furthermore, a comprehensive review of wormhole physics can be found in Ref. [19].
@@
Some authors have highlighted the potential existence of wormholes under extreme astrophysical conditions. Bhar et al. [28] investigated the possibility of sustaining static, spherically symmetric traversable wormhole geometries in GR, supported by dark energy and admitting conformal motion. Biswas et al. [29] proposed that wormholes could mimic supermassive black holes at galactic centers. Other extreme astrophysical environments where wormholes are hypothesized to exist include compact stars [30, 31, 32, 33, 34, 35, 36, 37, 38]. However, in Refs. [30, 31, 32, 33, 34, 35, 36, 37], the existence of hybrid systems composed of wormholes and compact stars is sustained by the presence of ghosts, which leads to undesirable consequences: both mixed star-plus-wormhole systems and pure wormholes are unstable [38]. Nojiri et al. [38] eliminated the ghosts by imposing the constraints given by the additional Lagrange multipliers.
@@
It has to be noted that although the model proposed in Ref. [38] has successfully constructed a formalism of ghost-free systems, it did not take into account the anisotropy that may arise. It is widely known that the matter inside compact stars like neutron stars (NSs) may be anisotropic, meaning that the radial and tangential pressures differ. Various factors can lead to such pressure anisotropy; including boson condensation, the presence of a solid core, different types of phase transitions, strong magnetic or electric fields, and the influence of modified spacetime geometry [39]. In line with these physical considerations, several studies on compact stars [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] have incorporated anisotropic pressure. Anisotropy plays an important role in the mass calculation of compact stars. For instance, Ref. [49] shows that anisotropy can enhance the mass of NSs, allowing them to reach the mass range of the secondary compact object in GW190814, which has been reported in Ref. [55] and hypothesized as the heaviest NS ever observed [56]. The review of anisotropic compact stars can be referred to Ref. [57].
@@
Another aspect that is not taken into account in the model presented in Ref. [38] is the presence of magnetic fields. The magnetic fields on NSs are typically in the range of $10^{12}$$10^{15}$[58]. There are classes of pulsars, such as anomalous X-ray pulsars (AXPs) and soft gamma repeater pulsars (SGRs), that have been identified as producing strong magnetic fields. SGRs are associated with supernova remnants, which correspond to young NSs [59]. Furthermore, observations of some AXPs also indicate that their surface magnetic fields are around $10^{14}-10^{15}$[59]. For further details on the observational diversity of magnetized NSs, see also Ref. [60], while a general review of NSs is provided in Ref. [61].
@@
One framework for describing magnetic fields in NSs is through chaotic magnetic fields, which has been investigated by Lopes & Menezes [62]. In their work, they also introduced an ansatz in which the magnetic field is coupled to the energy density of the star. An important advantage of adopting a chaotic magnetic field is the elimination of another anisotropy induced by the magnetic field, thereby simplifying the mathematical formulation. In addition to this simplification, the use of chaotic magnetic fields allows one to obtain a magnetic pressure that is independent of both direction and the choice of coordinate system, thus restoring the proper thermodynamic concept of pressure and ensuring consistency with field theory [62, 63]. Several studies have subsequently incorporated chaotic magnetic fields in their analyses [63, 64, 65, 66].
@@
Our aim in this work is to extend the wormhole plus neutron star (WH+NS) model constructed in Ref. [38] by incorporating the anisotropy of the neutron fluid and the effects of chaotic magnetic fields that may arise in such a system. Other distinctions of our work from previously studied WH+NS configurations constructed in Refs. [30, 31, 32, 33, 34, 35, 36, 37, 38] is the inclusion of the calculations of surface redshift and the echo time. Motivated by the possibility that WH+NS systems with an anisotropic neutron fluid and a chaotic magnetic field may contribute to gravitational-wave echoes characterized by the echo time, our results provide potential observational imprint for future LIGO/Virgo searches.
@@
The paper is organized as follows. In Sect. II, we introduce the formalisms of the anisotropic fluid and the chaotic magnetic field. Sec. III presents the formulation of GR coupled with two scalar fields, and demonstrates the elimination of ghosts. In this section, we also analyze the energy conditions and the travesability of the wormholes. In Sec. IV, we construct a model of non-magnetized WH+NS systems that consider anisotropy. Sec. V shows the model of anisotropic magnetized WH+NS systems. Note that, in both Secs. IV and V, we compare the obtained WH+NS systems with ordinary NSs; and also present the calculation results of the surface redshift and the echo time that may arise in such systems. Finally, Sec. VI is devoted to the conclusions of our study.
II Anisotropic fluid and chaotic magnetic field
@@
To make this section self-contained, in Subsect. II.1 we introduce the anisotropic fluid model employed in this study, while the formulation and ansatz of the chaotic magnetic field are given in Subsect. II.2.
II.1 Model of anisotropy
@@
It is important to consider the EMT of matter with anisotropic pressure, given by [67]
@@
where $\rho$ is the energy density, $p_{t}$ is the tangential pressure, and $\sigma=p_{r}-p_{t}$ denotes the anisotropy term, where $p_{r}$ is the radial pressure. At this stage, we have to emphasize that $\rho$, $p_{r}$, and $p_{t}$ belong to matter. In Eq. (1), $u^{\mu}$ is the four-velocity of the fluid, satisfying the normalization condition $u^{\mu}u_{\mu}=-1$, and $\xi^{\mu}$ is a unit radial four-vector, satisfying $\xi^{\mu}\xi_{\mu}=1$. By choosing $\xi_{\mu}=(0,\sqrt{g_{rr}},\>0,\>0)$, for static spherically symmetric configuration, we can easily find
@@
In this work, we consider Cosenza-Herrera-Esculpi-Witten (CHEW) model [68] in the realization of the WH+NS with anisotropic neutron fluid. The expression of the anisotropy term reads [45]
@@
II.2 Chaotic magnetic fields
@@
It should be noted that the standard treatment of magnetic fields in NSs, which leads to the appearance of anisotropy even though the original matter of the NSs is isotropic [70], complicates the mathematical formulation and prevents the stellar configuration from remaining spherically symmetric [71].
Another important point to note is that the exact mathematical profile of the magnetic field inside NSs remains unknown. Lopes & Menezes [62] proposed an ansatz in which the magnetic field is coupled to the matter energy density $\rho$, i.e.
@@
where $B_{s}$ denotes the magnetic field at the surface of the NS, $B_{0}$ represents the magnetic field expected at the core. For the WH+NS configuration, $B_{0}$ can be interpreted as the magnetic field expected at the wormhole’s throat. Moreover, $\rho_{0}$ denotes the central energy density of the non-magnetized NS configuration (without the wormhole) corresponding to the maximum mass, and $\eta$ is an arbitrary positive constant. It is worth noting that choosing $\eta$ to be very close to zero would make the magnetic field profile resemble a constant magnetic field. To simplify the calculation of the radius of the wormhole throat in Sec. V, we set $\eta=1$ throughout this work. Moreover, the magnetic field parameters that are varied in our analysis are both $B_{0}$ and $B_{s}$.
III GR coupled with two scalar fields and a magnetic field
For completeness, Subsect. III.1 contains the derivation of the field equations from the action, Subsect. III.2 addresses the procedure for eliminating the ghosts, and Subsect. III.3 analyzes the energy conditions and traversability of the wormholes.
III.1 Einstein field equations

### v3

@@ Abstract
In this paper, we formulate wormhole-plus-neutron-star (WH+NS) systems supported by two scalar fields, allowing for both pressure anisotropy of the neutron fluid and magnetic field. In general, such WH+NS systems contain ghosts; however, these ghosts can be eliminated. We find that the wormhole remains traversable regardless of whether anisotropy of the neutron fluid and/or magnetic fields are included. In particular, the null energy condition (NEC) remains violated in the vicinity of the wormhole throat, ensuring the traversable nature of the geometry. For magnetized configurations, the resulting WH+NS systems can become extremely massive, with ADM masses exceeding $8\,M_{\odot}$, and can exhibit large surface redshifts exceeding $z\simeq 1.5$. Furthermore, we analyze the gravitational-wave echo time of the systems, which serves as a potential observational imprint. Our results indicate that the echo time can vary depending on the fluid anisotropy and the magnetic field configuration, suggesting that WH+NS systems may provide distinctive signals of gravitational echo.
journal:
1 Introduction
@@
Wormholes are intriguing constructs that serve as tunnels connecting distant spacetime points Dai2020 ; Turimov2025 ; Bronnikov2023 ; Cataldo2017 ; Yousaf2025 ; Prat2015 ; Dai2019 ; Godani2019 . Both distinct spacetime domains are linked through a throat-like geometry. Originally, wormholes emerged as exact solutions of the Einstein field equation Zangeneh2025 . Due to their geometric nature, these exotic objects may potentially exist in curved spacetimes within a broad class of gravity theories Bambi2021 . Interestingly, wormholes have been associated with rapid interstellar travel, warp drives Shinkai2015 , and candidates for active galactic nuclei Piotrovich .
@@
Historically, the first insight into what would later be interpreted as a wormhole can be traced back to Flamm’s 1916 work Flamm1916 (see Ref. Flamm2015 for the English translation), where he analyzed the Schwarzschild spacetime. In 1935, Einstein & Rosen Einstein1935 found a solution bearing a resemblance to a wormhole, which was identified as a “bridge” that connects two identical physical spaces. This work gave rise to the term “Einstein–Rosen bridge”. The term “wormhole” was later introduced in 1957 by Misner & Wheeler Misner1957 . It is worth noting that the putative wormhole in the Einstein–Rosen bridge is not traversable, since it is colocated with the blackhole’s singularity. The surface of the blackhole is the event horizon, which cannot be a wormhole Radhakrishnan2024 .
@@
The formalism of traversable wormholes was introduced by Morris & Thorne Morris1988 in their seminal 1988 paper. In constructing traversable wormholes, the null energy condition (NEC) must be violated by the energy-momentum tensor (EMT) Lobo2016 . It should be noted that there are three other energy conditions: the weak energy condition (WEC), the strong energy condition (SEC), and the dominant energy condition (DEC) Radhakrishnan2024 . The violation of NEC implies that all the energy conditions are not validated Moraes2019 . In general relativity (GR), the violation of NEC typically requires the presence of exotic matter Kokubu2020 . Many authors Chianese2017 ; Mustafa2020 ; AzregAinou2015 ; Kuhfittig2006 ; Bhattacharya2019 ; Sreekumar2025 have included the exotic matter in their studies on wormholes. Furthermore, a comprehensive review of wormhole physics can be found in Ref. Lobo2016 . Refs. Falco1 ; Falco2 ; Falco3 ; Falco4 ; Falco5 ; Falco6 work on the development of possible astrophysical techniques to detect the wormholes.
@@
Some authors have highlighted the potential existence of wormholes under extreme astrophysical conditions. Bhar et al. Bhar2016 investigated the possibility of sustaining static, spherically symmetric traversable wormhole geometries in GR, supported by dark energy and admitting conformal motion. Biswas et al. Biswas2024 proposed that wormholes could mimic supermassive black holes at galactic centers. Refs. Hao1 ; Hao2 work on wormhole-plus-black-hole systems. Other extreme astrophysical environments where wormholes are hypothesized to exist include compact stars Dzhunushaliev2011 ; Dzhunushaliev2012 ; Dzhunushaliev2013 ; Dzhunushaliev2014 ; Aringazin2015 ; Dzhunushaliev2015 ; Dzhunushaliev2016 ; Dzhunushaliev2023 ; Nojiri2024 . However, in Refs. Dzhunushaliev2011 ; Dzhunushaliev2012 ; Dzhunushaliev2013 ; Dzhunushaliev2014 ; Aringazin2015 ; Dzhunushaliev2015 ; Dzhunushaliev2016 ; Dzhunushaliev2023 , the existence of hybrid systems composed of wormholes and compact stars is sustained by the presence of ghosts, which leads to undesirable consequences: both mixed star-plus-wormhole systems and pure wormholes are unstable Nojiri2024 . Nojiri et al. Nojiri2024 eliminated the ghosts by imposing the constraints given by the additional Lagrange multipliers.
@@
It has to be noted that although the model proposed in Ref. Nojiri2024 has successfully constructed a formalism of ghost-free systems, it did not take into account the anisotropy that may arise. It is widely known that the matter inside compact stars like neutron stars (NSs) may be anisotropic, meaning that the radial and tangential pressures differ. Various factors can lead to such pressure anisotropy; including boson condensation, the presence of a solid core, different types of phase transitions, strong magnetic or electric fields, and the influence of modified spacetime geometry Pattersons2021 . In line with these physical considerations, several studies on compact stars Pattersons2021 ; Sulaksono2014 ; Sulaksono2015 ; Silva2015 ; Setiawan2017 ; Lopes2019 ; Rizaldy2024 ; Posada2024 ; Becerra2024 ; Pattersons2025 ; Yusmantoro2025 ; Sakti2016 ; Becerra2024b ; Becerra2025 ; Torres2019 ; Rahmansyah2020 have incorporated anisotropic pressure. Anisotropy plays an important role in the mass calculation of compact stars. For instance, Ref. Pattersons2025 shows that anisotropy can enhance the mass of NSs, allowing them to reach the mass range of the secondary compact object in GW190814, which has been reported in Ref. Abbott2019 and hypothesized as the heaviest NS ever observed Fattoyev2020 . The review of anisotropic compact stars can be referred to Ref. Kumar2022 .
@@
Another aspect that is not taken into account in the model presented in Ref. Nojiri2024 is the presence of magnetic fields. The magnetic fields on NSs are typically in the range of $10^{12}$$10^{15}$Lander2009 . There are classes of pulsars, such as anomalous X-ray pulsars (AXPs) and soft gamma repeater pulsars (SGRs), that have been identified as producing strong magnetic fields. SGRs are associated with supernova remnants, which correspond to young NSs Rizaldy2018a . Furthermore, observations of some AXPs also indicate that their surface magnetic fields are around $10^{14}-10^{15}$Rizaldy2018a . For further details on the observational diversity of magnetized NSs, see also Ref. Enoto2019 , while a general review of NSs is provided in Ref. Vidaña2018 .
@@
One framework for describing magnetic fields in NSs is through chaotic magnetic fields, which has been investigated by Lopes & Menezes Lopes2015 . In their work, they also introduced an ansatz in which the magnetic field is coupled to the energy density of the star. An important advantage of adopting a chaotic magnetic field is the elimination of another anisotropy induced by the magnetic field, thereby simplifying the mathematical formulation. In addition to this simplification, the use of chaotic magnetic fields allows one to obtain a magnetic pressure that is independent of both direction and the choice of coordinate system, thus restoring the proper thermodynamic concept of pressure and ensuring consistency with field theory Lopes2015 ; Pattersons2024a . Several studies have subsequently incorporated chaotic magnetic fields in their analyses Pattersons2024a ; Wu2017 ; Lopes2020 ; Pattersons2024b .
@@
Our aim in this work is to extend the wormhole plus neutron star (WH+NS) model constructed in Ref. Nojiri2024 by incorporating the anisotropy of the neutron fluid and the effects of chaotic magnetic fields that may arise in such a system. Other distinctions of our work from previously studied WH+NS configurations constructed in Refs. Dzhunushaliev2011 ; Dzhunushaliev2012 ; Dzhunushaliev2013 ; Dzhunushaliev2014 ; Aringazin2015 ; Dzhunushaliev2015 ; Dzhunushaliev2016 ; Dzhunushaliev2023 ; Nojiri2024 is the inclusion of the calculations of surface redshift and the echo time. Motivated by the possibility that WH+NS systems with an anisotropic neutron fluid and a chaotic magnetic field may contribute to gravitational-wave echoes characterized by the echo time, our results provide potential observational imprint for future LIGO/Virgo searches.
We emphasize that the presence of a magnetic field should not be regarded as a merely trivial extension of the system studied in Ref. Nojiri2024 . This is clearly reflected in the fact that the magnetized WH+NS configuration admits a different analytical solution for the throat radius. In addition, we obtain potential observational imprints of magnetized WH+NS systems that have not been explored previously, most notably their gravitational mass, which can exceed $8\,M_{\odot}$, being noticeably higher than that of the ordinary WH+NS systems investigated in Ref. Nojiri2024 , and their surface redshift, which can be larger than 1.5—both significantly higher than those of ordinary NSs. Moreover, we find that the combined effects of fluid anisotropy and the magnetic field lead to distinctive echo times, providing further observationally relevant signatures.
@@
The paper is organized as follows. In Sect. 2, we introduce the formalisms of the anisotropic fluid and the chaotic magnetic field. Sec. 3 presents the formulation of GR coupled with two scalar fields, and demonstrates the elimination of ghosts. In this section, we also analyze the energy conditions and the travesability of the wormholes. In Sec. 4, we construct a model of non-magnetized WH+NS systems that consider anisotropy. Sec. 5 shows the model of anisotropic magnetized WH+NS systems. Note that, in both Secs. 4 and 5, we compare the obtained WH+NS systems with ordinary NSs; and also present the calculation results of the surface redshift and the echo time that may arise in such systems. Finally, Sec. 6 is devoted to the conclusions of our study.
2 Anisotropic fluid and chaotic magnetic field
@@
To make this section self-contained, in Subsect. 2.1 we introduce the anisotropic fluid model employed in this study, while the formulation and ansatz of the chaotic magnetic field are given in Subsect. 2.2.
2.1 Model of anisotropy
@@
It is important to consider the EMT of matter with anisotropic pressure, given by Pretel2023
@@
where $\rho$ is the energy density, $p_{t}$ is the tangential pressure, and $\sigma=p_{r}-p_{t}$ denotes the anisotropy term, where $p_{r}$ is the radial pressure. At this stage, we have to emphasize that $\rho$, $p_{r}$, and $p_{t}$ belong to matter. In Eq. (1), $u^{\mu}$ is the four-velocity of the fluid, satisfying the normalization condition $u^{\mu}u_{\mu}=-1$, and $\xi^{\mu}$ is a unit radial four-vector, satisfying $\xi^{\mu}\xi_{\mu}=1$. The vector $\xi_{\mu}$ is physically an auxiliary quantity that parametrizes the anisotropic stresses in the neutron fluid. Its components are fully determined by the geometry and matter profiles, in such a way so that $p_{r}=\xi^{\mu}\xi^{\nu}T_{\text{matter}\>\mu\nu}$, and $p_{t}=\varsigma^{\mu}\varsigma^{\nu}T_{\text{matter}\>\mu\nu}$, where $\varsigma^{\mu}$ is a unit spacelike vector (orthogonal to $u^{\mu}$ and $\xi^{\mu}$anisoadd .
By choosing $\xi_{\mu}=(0,\sqrt{g_{rr}},\>0,\>0)$, for static spherically symmetric configuration, we can easily find
@@
In this work, we consider Cosenza-Herrera-Esculpi-Witten (CHEW) model Cosenza1981 in the realization of the WH+NS with anisotropic neutron fluid.
For the static spherically symmetric spacetime, the metric describing the interior of ordinary NSs is given by
${\color[rgb]{0,0,0}ds^{2}=-e^{2\nu(r)}dt^{2}+e^{2\lambda(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),}$ (3)
where $\lambda(r)=\left(1-\frac{2m(r)}{r}\right)^{-1}$, and $m(r)$ denotes mass of the NSs as a function of $r$. Using the EMT shown in Eq. (1), Tolman-Oppenheimer-Volkoff (TOV) equation for anisotropic NSs reads
${\color[rgb]{0,0,0}\frac{dp_{r}}{dr}=-\frac{d\nu}{dr}(\rho+p_{r})-\frac{2\sigma}{r}.}$ (4)
Authors in Ref. Cosenza1981 considered that $\sigma/r$ can be written as
${\color[rgb]{0,0,0}\frac{\sigma}{r}=f(p_{r},r)(\rho+p_{r}),}$ (5)
where $f(p_{r},r)$ is an arbitrary function. Motivated to obtain a minimal modification of the TOV equation, authors in Ref. Cosenza1981 assumed $f(p_{r},r)=\frac{1-h}{2}\frac{d\nu}{dr}$, where $h$ is a constant. On the other words, it is not assumed for any specific physical reason but only because it transforms TOV equation into the simple form, i.e.
${\color[rgb]{0,0,0}\frac{dp_{r}}{dr}=-h\frac{d\nu}{dr}(\rho+p_{r})}.$ (6)
Finally, the expression of the anisotropy term can also be written as Rahmansyah2020
@@
2.2 Chaotic magnetic fields
@@
It should be noted that the standard treatment of magnetic fields in NSs, which leads to the appearance of anisotropy even though the original matter of the NSs is isotropic Bordbar2022 , complicates the mathematical formulation and prevents the stellar configuration from remaining spherically symmetric Mallick2014 .
According to Ref. Lopes2015 , the chaotic magnetic field approximation adopted here is physically motivated by the expectation that magnetic fields generated during the turbulent core-collapse supernova phase are highly disoriented. In this framework, the magnetic pressure is assumed to be statistically isotropic, so that no global preferred direction exists and anisotropic magnetic stresses associated with ordered field configurations are effectively averaged out. This allows the magnetic contribution to be incorporated consistently within the spherically symmetric formalism through an effective isotropic pressure. A limitation of the model is that it is expected that the magnetic field evolves and becomes at least partially oriented. The oriented magnetic field is a necessity to the theory of pulsars as celestial lighthouse. In this work, the the chaotic magnetic field is employed as an approximation to calculate the influence of the magnetic field in NS properties within the spherically symmetric TOV formalism.
Another important point to note is that the exact mathematical profile of the magnetic field inside NSs remains unknown. Lopes & Menezes Lopes2015 proposed an ansatz in which the magnetic field is coupled to the matter energy density $\rho$, i.e.
@@
where $B_{s}$ denotes the magnetic field at the surface of the NS, $B_{0}$ represents the magnetic field expected at the core. For the WH+NS configuration, $B_{0}$ can be interpreted as the magnetic field expected at the wormhole’s throat. Moreover, $\rho_{0}$ denotes the central energy density of the non-magnetized NS configuration (without the wormhole) corresponding to the maximum mass, and $\eta$ is an arbitrary positive constant. It is worth noting that choosing $\eta$ to be very close to zero would make the magnetic field profile resemble a constant magnetic field. To simplify the calculation of the radius of the wormhole throat in Sec. 5, we set $\eta=1$ throughout this work. Moreover, the magnetic field parameters that are varied in our analysis are both $B_{0}$ and $B_{s}$.
It is important to note that the internal magnetic field profile of NSs is still subject to significant uncertainties. For this reason, it is customary to introduce phenomenological but physically motivated ansatz for the magnetic field profile. Different prescriptions have been proposed in the literature, including magnetic field profiles coupled to the baryon number density, as in Ref. Mallick2014 , as well as profiles coupled to the energy density.
In the latter approach, i.e. the ansatz proposed in Ref. Lopes2015 , the energy density can enter directly into the TOV equations governing the macroscopic structure of ordinary NSs. The parameter $\rho_{0}$ serves as a reference energy density scale, while the parameter $\eta$ controls the strength of the magnetic field. This choice provides a convenient and physically motivated parametrization for exploring the impact of strong magnetic fields on NS properties.
3 GR coupled with two scalar fields and a magnetic field
For completeness, Subsect. 3.1 contains the derivation of the field equations from the action, Subsect. 3.2 addresses the procedure for eliminating the ghosts, and Subsect. 3.3 analyzes the energy conditions and traversability of the wormholes.
3.1 Einstein field equations